the first and last teem of an AP are 0 and 108.if the sum the series is 702.find d number of term in the AP and the common difference
Answer by Boreal(15235)

Sum=(n/2)(a1+an)
702=(n/2)(0+108)
1404=108n
n=13 terms
last term is a1+d(n-1)
108=d(12)
d=9 difference
the series is 0,9,18,27,36,45,54,63,72,81,90,99,108
that adds to 702.